1 2 3 4 5 6 7 8 9  1 2 3 4 5 6 7 8 9 1 2 3 4 5 6 7 8 9 1 2 3 4 5 6 7 8 9

This frame size is 125.2mm × 199.2mm. When printed on B5E size paper, the frame should be one inch from each edge of the paper, where one inch is equal to 25.4mm. Please control the top lefthand corner of this frame should be one inch from the left and top edges of the paper. The ticks of the rulers at the four edges are 0.1 inch apart.